(5.2.27) Based on a​ poll, 60​% of adults believe in reincarnation. Assume that 6 adults are randomly​ selected, and find the indicated probability.
A. The probability that exactly 5 of the 6 adults believe in reincarnation is ?
1
Expert's answer
2020-12-02T17:22:37-0500

P = 60 % = 0.60

n = 6

Evaluate P(x) if x = 5

"P(x = 5) = \frac{6!}{5!(6-5)!} \times 0.6^5 \times (1 – 0.6)^{6-5} = 6 \times 0.6^5 \times 0.4^{1} = 0.1866"

Answer: 0.1866 or 18.66 %